40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 14 01:56:20 2013, Jenne, Update, Green Locking, Xend Green tweaked 
    Reply  Tue May 14 02:19:13 2013, Koji, Update, Green Locking, Xend Green tweaked 
    Reply  Tue May 21 18:18:23 2013, Manasa, Update, Green Locking, Xend Green aligned 
       Reply  Tue May 21 23:29:57 2013, Manasa, Update, Green Locking, Xend Green aligned 
          Reply  Tue May 28 19:35:17 2013, Manasa, Update, Green Locking, Still searching for the X green beat note 
             Reply  Thu May 30 01:02:41 2013, Manasa, Update, Green Locking, Found it! IMG_0598.JPGIMG_0599.JPG
                Reply  Mon Jun 3 21:09:55 2013, Manasa, Update, Green Locking, X green beatnote through beatbox IMG_0608.JPG
                   Reply  Tue Jun 11 22:16:56 2013, Manasa, Summary, Green Locking, ALS for X arm 
                      Reply  Thu Jun 13 22:31:12 2013, Manasa, Summary, Green Locking, X arm ALS IMG_0659.JPGALS_stabilized_13June2013.png
                         Reply  Fri Jun 14 07:06:19 2013, Koji, Summary, Green Locking, X arm ALS 
Message ID: 8653     Entry time: Thu May 30 01:02:41 2013     In reply to: 8651     Reply to this: 8666
Author: Manasa 
Type: Update 
Category: Green Locking 
Subject: Found it! 

X green beat note found!

Key points
1. Near-field and far-field alignment on the PSL table. The near-field alignment checked by looking at the camera and the far-field alignment checked by allowing the beams to propagate by removing the DC PD.
2. Check laser temperature and get a sense of how the offset translates to the actual laser temperature.
3. Get an idea of the expected temperature of laser using the plot in elog.

Data
PSL laser temperature = 31.45 deg C

X end laser temperature = 39.24 deg C
C1-ALS-X_SLOW_SEERVO2_OFFSET = 4810
Amplitude of beat note = -40dBm

I do not understand why
1. The amplitude of beatnote falls linearly with frequency (peak traced using 'hold' option of the spectrum analyzer).
2. I found the beat note at the RF output of the PD. Earlier, while I was trying to search for the beatnote from the RFmon output of the betabox, there was a strong peak at 29.6MHz that existed even when the green shutters were closed. It's source has to be traced.

Next
Solve beatbox puzzle and lock arm using ALS.

IMG_0598.JPGIMG_0599.JPG

ELOG V3.1.3-